through:(4,-4), parallel to y =-x-4
Answer by Alan3354(69443)

Do it like this one.
======================
through:(-3,-3), parallel to y =7/3x+3
-----------
The slope is the same, m = 7/3
Parallel lines have the same slope.
-------
Use y = mx + b and the point to find b, the y-intercept
-3 = (7/3)*-3 + b
b = 4
--------
--> y = (7/3)x + 4